A quality control want to estimate the proportion of high quality goods out of a batch of products with a 90% confidence level and a margin of error of 5%. How may products must he tes

1
Expert's answer
2022-06-10T14:22:29-0400

A 90% confidence interval has a z-score (a critical value) "z_c=1.6449."



"ME=z_c\times\sqrt{\dfrac{p(1-p)}{n}}""=1.6449\times\sqrt{\dfrac{0.5(1-0.5)}{n}}=0.05""n=(16.449)^2""n=271"